A particle moves over the sides of an equilateral triangle of side l with constant speed v along the sides A, B and C.Then
1.What is the magnitude of average velocity from A to C?
2.What is the magnitude of average acceleration from A to C?

Grade:10

1 Answers

Arun
5 years ago
Given :
Average velocity= Total displacement/total time
Total displacement from A to C will be 2l
V=2l/T--------(1)
T=2l/V-----(2)
acceleration =change in velocity/time=V/T=V/2l/V----(from equation 2)
=V²/2l
∴The magnitude of average velocity from A to C is 2l/T 
The magnitude of average acceleration from A to C is V²/2l